OpenStudy (anonymous):

solve the triangle ABC in which A=66, B=42 and ac=12cm

OpenStudy (anonymous):

What type of triangle is it?

OpenStudy (wolf1728):

tan (66°) = hgt/x hgt = tan (66°) * x tan(72°) = hgt/(12-x) hgt = tan(72°) * (12-x) tan(66°) * x = tan(72°) * (12-x) 2.2460367739 * x = 3.0776835372 * (12-x) 2.2460367739 * x = 36.9322024461 -3.0776835372*x 5.3237203111 *x = 36.9322024461 x = 6.9372920229 tan(66°) = hgt/x hgt = 2.2460367739 * 6.9372920229 hgt = 15.5814129947 AB² = x² + hgt² AB² = 48.126020611 + 242.7804309114 AB² = 290.9064515224 AB = 17.0559799344 BC² = hgt² + (12-x)² BC² = 242.7804309114 + 25.6310120614 BC² = 268.4114429728 BC = 16.3832671642